Help please!!!! I don't understand the process of getting the answer.
The following exercise demonstrates applications of linear functions. Analyze the information given to determine the equation of the relevant line or lines. Use the linear equation to determine the requested information.
The Cookie Store sells cookies for 44 cents each. The store has monthly fixed costs of $425 per month and unit costs of 22 cents per cookie. How many cookies must it sell each month to make a profit? (Round answer up to the nearest whole number.)

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

The Cookie Store sells cookies for 44 cents each.
---------
The store has monthly fixed costs of $425 per month and unit costs of 22 cents per cookie. How many cookies must it sell each month to make a profit? (Round answer up to the nearest whole number.)
----
Revenue = 44x where x is # of cookies sold.
----
Cost = 42500 + 22x where x is # of cookies made.
-----
Profit = Revenue - Cost = 44x -42500-22x = 22x -42500
----
To make a Profit 22x-42500 > 0
22x > 42500
x >= 1932 cookies must be sold per month.
==============================================
cheers,
Stan H.

Answer by mananth(16949)

Fixed cost = $425 / month
Cookies sell at $0.42
Cost = $0.22
profit per cookie = 0.22
Let number sold per month be x
C(x) = 425+0.22x
S(x)= 0.42x
To make a profit
Sales> Cost
0.42x>425+0.22x
0.42x-0.22x>425
0.2x>425
x>425/0.2
x>2125